OpenStudy (anonymous):

Hello all, Would someone kindly point me in the right direction for the following problem. Two Hydrogen atoms oscillate symmetrically relative to their centre mass. Force between the two is: \[F(x)=\frac{ 2U }{ b }(e ^{-2(x-x0)/b} - e ^{(x-x0)/b})\] Motion of the atom is simple harmonic, thus the force between the two atoms behaves likes a spring and the two atoms act like two masses joined by a spring. The frequency of oscillation is 1.31 x 10^14, mass of each atom is 1.67 x 10^-27. Find the value of the spring constant.

OpenStudy (amriju):

could u plz write the expression more clearly?

OpenStudy (fifciol):

\[m \ddot x+F(x)=0\] But it is said that motion is simple harmonic, if this is a case you can calculate k using this equation; \[\omega=\sqrt{\frac{ k }{m }}=2\pi f \Rightarrow k=4\pi^2f^2m\]

OpenStudy (fifciol):

I tried to solve that DE with that weird F(x) function but even wolfram can't do it so I think this simple harmonic solution is correct

OpenStudy (anonymous):

Thanks so much Fifciol. Apologies if the F(x) expression was not clear. The correct one is: \[f(x) = \frac{ 2U }{ b } (e^{-2(x-x _{0})/b} - e ^{(x-x _{0})/b})\]

OpenStudy (anonymous):

The simple harmonic solution looks correct but I feel im missing the point of the question. I cant solve that DE and am stuck at that point....